OpenStudy (anonymous):

r + (-12) = 19 can someone please help me i dont understand this problem

OpenStudy (anonymous):

help me after

OpenStudy (anonymous):

you can rewrite it first then undo

OpenStudy (anonymous):

r - 12 = 19 now undo -12 by adding 12 to both sides

OpenStudy (anonymous):

so r = 19+12 then plug value back in original equation to check

OpenStudy (anonymous):

p + (-9) = -9 now what about this problem can you help me with it please

OpenStudy (anonymous):

adding a negative is same as subtracting a positive

OpenStudy (anonymous):

same concept change to p-9=-9

OpenStudy (anonymous):

undo -9 by adding 9 to both sides to keep the equation "balanced"-whatever you do to one side, you must do to the other

OpenStudy (anonymous):

so what would r = ?

OpenStudy (anonymous):

9 i think

OpenStudy (anonymous):

well what is -9 + 9

OpenStudy (anonymous):

r = 31

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

oh ok i am a little slow so

OpenStudy (anonymous):

they are opposites...when you add a number and its opposite you get 0

OpenStudy (anonymous):

r definitely equals 31.

OpenStudy (anonymous):

its ok...math takes time ; )

OpenStudy (anonymous):

p=0

OpenStudy (anonymous):

yes for original problem : ) we are working another

OpenStudy (anonymous):

oh i see.

OpenStudy (anonymous):

I'm trying to show process then answer so chill ; )

OpenStudy (anonymous):

Have at it.

OpenStudy (anonymous):

thanks ; )

OpenStudy (anonymous):

p + (-9) = -9 p-9 = -9 + 9 +9 p=0 check: 0 + (-9) = -9 -9=-9